OpenStudy (anonymous):

Find dy/dx by implicit differentiation. y^2+xy=x^2+1

OpenStudy (michele_laino):

Please you have to compute the first derivative of both sides of your equation, using the chain-rule

OpenStudy (anonymous):

I don't understand the chain rule.

OpenStudy (anonymous):

How would I derive xy

OpenStudy (michele_laino):

for example, I apply the cahin-rule as follows: \[\frac{d}{{dx}}\left( {xy} \right) = y + xy'\]

OpenStudy (anonymous):

okay. would the 2nd part be just x?

OpenStudy (michele_laino):

no, since you have y=y(x) and please keep in mind you are computing the first derivative with respect to the x-variable

OpenStudy (michele_laino):

hint: \[\frac{d}{{dx}}\left( {{y^2}} \right) = 2yy'\]

OpenStudy (anonymous):

Thank you for being kind to me, I'm so hopelessly lost.

OpenStudy (michele_laino):

that is the chain-rule

OpenStudy (anonymous):

My textbook doesn't even mention that. So it is second derivative?

OpenStudy (michele_laino):

no it is the first derivative, it is the first deivative of a function of function. namely: \[\frac{{d{y^2}}}{{dx}} = \frac{{d{y^2}}}{{dy}} \cdot \frac{{dy}}{{dx}} = 2y \cdot y'\]

OpenStudy (michele_laino):

please you have to search for first derivative of function of function, or first derivative of composite functions, on your textbook

OpenStudy (anonymous):

according to my textbook a first derivative function is: \[y ^{1}=\frac{ dy }{dx }=f \prime \left( x \right)\]

OpenStudy (anonymous):

how would I incorporate that into the question?

OpenStudy (michele_laino):

yes! that is the fist derivative of y=f(x), nevertheless kep in mind that yo have this function: \[g\left( x \right) = y\left( x \right) \cdot y\left( x \right)\] Next I apply the Leibniz rule, and I get: \[\frac{d}{{dx}}g\left( x \right) = \frac{{dy}}{{dx}} \cdot y + y \cdot \frac{{dy}}{{dx}} = 2y \cdot \frac{{dy}}{{dx}}\]

OpenStudy (michele_laino):

oops.. nevertheless keep in mind...

OpenStudy (anonymous):

I do not believe we covered the Leibniz rule

OpenStudy (anonymous):

Where does the 2y come from?

OpenStudy (michele_laino):

please note that sometime the "Leibniz rule" is called "product rule"

OpenStudy (anonymous):

is that the y=uv?

OpenStudy (michele_laino):

there is 2y since I have added similar terms, namely: \[\frac{{dy}}{{dx}} \cdot y + y \cdot \frac{{dy}}{{dx}} = y \cdot \frac{{dy}}{{dx}} + y \cdot \frac{{dy}}{{dx}}\] since: \[\frac{{dy}}{{dx}} \cdot y = y \cdot \frac{{dy}}{{dx}}\]

OpenStudy (michele_laino):

yes, it is the same rule that we have to apply in order to find the derivative of y=u*v

OpenStudy (anonymous):

okay that makes sense. Since you've applied both rules to the left side of the equation. What do I do with the right side?

OpenStudy (michele_laino):

the right side is very easy, please you have to apply this identity: \[\frac{d}{{dx}}{x^n} = n{x^{n - 1}}\]

OpenStudy (michele_laino):

furthermore, please keep in mind that, we have: \[\frac{d}{{dx}}1 = 0\]

OpenStudy (anonymous):

okay so for the end of this problem do I come up with a number for an answer or another equation?

OpenStudy (anonymous):

Because we aren't using real numbers necessarily, just applying rules. Do these rules satisfy the equation?

OpenStudy (michele_laino):

yes! all theorems of differential calculus can be applied to all problems which involve real functions of real variable, namely y=f(x)

OpenStudy (michele_laino):

of course we can apply those theorems also in order to solve problems with implicit functions

OpenStudy (anonymous):

and the equation I have is implicit.

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

Okay so first I have to solve for the left side by using the chain rule then the product rule? And for the right side you said to use the power rule?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

are there any other rules I need to apply?

OpenStudy (michele_laino):

another rule that you can have to apply is this: \[\frac{d}{{dx}}\left( {f + g} \right) = \frac{d}{{dx}}f + \frac{d}{{dx}}g\]

OpenStudy (anonymous):

but I do not have g? How do I apply this rule?

OpenStudy (michele_laino):

for example at the left side you have these function: \[\begin{gathered} f = {y^2} \hfill \\ g = xy \hfill \\ \end{gathered} \] please note that the symbols f and g are used in a general sense here

OpenStudy (anonymous):

OH okay! That clears it up

OpenStudy (michele_laino):

if you like I can rewrite the above identity as follows: \[\frac{d}{{dx}}\left( {P + Q} \right) = \frac{d}{{dx}}P + \frac{d}{{dx}}Q\]

OpenStudy (michele_laino):

and: \[\begin{gathered} P\left( x \right) = {y^2} \hfill \\ Q\left( x \right) = xy \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

\[\frac{ d }{ dx }(y^2+xy)=\frac{ d }{ dx }(y^2)+\frac{ d }{ dx }(xy)\] is that what it should look like?

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

how would I go about to simplify that further? Is this the part where I apply the chain rule?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

okay now I've solved it. Thank you so much for your patience!

OpenStudy (michele_laino):

Thank you!

OpenStudy (anonymous):

what is the volume in cubic inches of a cube whose total surface area is 96 square inches?

OpenStudy (anonymous):

can you help me out about the question?

OpenStudy (phi):

you should make a new post for your new question. but to answer it, notice that a cube has 6 (all equal ) sides, so one side has area 96/6 = 16 square inches. now you can figure out the length of 1 side of the square. the volume will be the side to the 3rd power